OpenStudy (anonymous):

If a+b=3(c+d) what is the average of a, b,c, and d?

OpenStudy (nikita2):

(a+b+c+d)/4 = (3(c+d) + c + d)/4 = 4(c+d)/4 = c+d

OpenStudy (anonymous):

Thank You!!:D